OpenStudy (anonymous):

dx/dy for x^5 = -xy^3 at (1,-1)

OpenStudy (anonymous):

again?

OpenStudy (anonymous):

yeah mate i got a homework load i gotta have some of these to study on. i gotta be ready

OpenStudy (anonymous):

isn't this the same one as before?

OpenStudy (anonymous):

if my memory serves me it is \(-\frac{4}{3}\)

OpenStudy (anonymous):

We used implicit differentiation on that one. For this, we have to find dx/dy

OpenStudy (anonymous):

oooh sorry

OpenStudy (anonymous):

Yeah, plus we got (1,-1) the points for this one

OpenStudy (anonymous):

flip it

OpenStudy (anonymous):

xy^3?

OpenStudy (anonymous):

take your answer to the last one' it was \[-\frac{5x^4+y^3}{3xy^2}\] and flip it

OpenStudy (anonymous):

3xy^2/5x^4+y^3

OpenStudy (anonymous):

im a pathetic case i know

OpenStudy (anonymous):

yaeh that is right, cept you dropped the minus sign

OpenStudy (anonymous):

so, -3xy^2/5x^4+y^3, right mate?

OpenStudy (anonymous):

\[\frac{dy}{dx}=-\frac{5x^4+y^3}{3xy^2}\] \[\frac{dx}{dy}=-\frac{3y^2}{5x^4+y^3}\] almost as if \(\frac{dx}{dy}\) and \(\frac{dy}{dx}\) were actual fractions hence the notation

OpenStudy (anonymous):

oh that's the answer, ya?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

you can start from scratch if you like, but that is silly after you did all the work now at \((1,-1)\) we got \(-\frac{4}{3}\) the first time this time you will get ... \(-\frac{3}{4}\)

OpenStudy (anonymous):

Wait, so is the answer -3/4 or the other one? gahh...

OpenStudy (anonymous):

\(\frac{dy}{dx}\) at \((1,-1)\) was \(-\frac{4}{3}\) and so \(\frac{dx}{dy}\) and \((1,-1)\) will be \(-\frac{3}{4}\)

OpenStudy (anonymous):

cheers mate love ya always a helpin' hand

OpenStudy (anonymous):

cheers to you